\newpage
{\footnotesize
\begin{thebibliography}{00}

\bibitem{beauchemin1988} P. Beauchemin, G. Brassard, A Generalization of Hellman's Extension to Shannon's Approach to Cryptography, {\it J. Cryptology}, {\bf 1} (1988), 129-131.
\bibitem{cover2006} T. M. Cover, J. A. Thomas, {\it Elements of Information Theory}, 2nd ed., Wiley-Interscience, 2006.
\bibitem{diaconis2009} P. Diaconis, The Markov Chain Monte Carlo Revolution, {\it Bulletin of the American Mathematical Society}, {\bf 46} (2009), 179-205.
\bibitem{fathi2016} B. Fathi-Vajargah, M. Kanafchian, Improved Markov Chain Monte Carlo Method for Cryptanalysis of Substitution-Transposition Cipher, {\it Monte Carlo Methods and Applications}, {\bf 22} (2016), 323-335.
\bibitem{hellman1977} M. Hellman, An Extension of Shannon's Theory Approach to Cryptography, {\it IEEE Trans. Inf. Theory}, {\bf 23} (1977), 289-294.
\bibitem{shannon1948} C. E. Shannon, A Mathematical Theory of Communication, {\it Bell Syst. Tech. J.}, {\bf 27} (1948), 379-423.
\bibitem{shannon1949} C. E. Shannon, Communication Theory of Secrecy Systems, {\it Bell Syst. Tech. J.}, {\bf 28} (1949), 656-715.
\bibitem{shannon1951} C. E. Shannon, Prediction and Entropy of Printed English, {\it Bell Syst. Tech. J.}, {\bf 30} (1951), 50-64.

\end{thebibliography}}

\documentclass[12pt]{article}
\usepackage{amsmath,amssymb,amstext,dsfont,fancyvrb,float,fontenc,graphicx,subfigure,theorem,hyperref}
\usepackage[utf8]{inputenc}

\usepackage[letterpaper]{geometry}
\usepackage[strict]{changepage}

\def\abstractname{Abstract -}   
\def\abstract{\begin{adjustwidth}{1cm}{1cm} \par    \footnotesize \noindent {\bf \abstractname} 
\def\endabstract{ \end{adjustwidth} \smallskip }}

{\theorembodyfont{\itshape}\newtheorem{theorem}{Theorem}[section]}
{\theorembodyfont{\itshape}}
{\theorembodyfont{\itshape}}
{\theorembodyfont{\itshape}}
{\theorembodyfont{\itshape}}
{\theorembodyfont{\rm}}
{\theorembodyfont{\rm}\newtheorem{remark}[theorem]{Remark}}

\title{\Large\bf Revisiting the Unicity Distance through a Channel Transmission Perspective}
\author{\sc F. Lin}

\begin{document}
\setcounter{page}{1}
\maketitle

\vskip 1.5em

\begin{abstract}
This paper revisits the classical notion of unicity distance from an enlightening perspective grounded in information theory, specifically by framing the encryption process as a noisy transmission channel. Using results from reliable communication theory, we derive a simple information-theoretic proof of the same unicity distance formula as in Shannon's classical result and a channel transmission interpretation of the unicity distance.
\end{abstract}

\begin{keywords}
unicity distance; reliable communication; channel coding theorem
\end{keywords}

\section{Introduction}
\label{sec:Introduction}
The unicity distance represents the amount of ciphertext required, on average, for an attacker to uniquely determine the encryption key.
In cryptography, encryption is a process that transforms a message, called plaintext, into a ciphertext. Encryption is represented as a function \( E \) that maps the plaintext \( P \) and a key \( K \) to the ciphertext \( C \).
\[
C = E_K(P)
\]

We consider only symmetric encryption scheme, so decryption is the reverse process, represented by a decryption function \( D \), which takes the ciphertext \( C \) and the key \( K \) to recover the original plaintext:

\[
P = D_K(C).
\]

In this paper, we revisit the classical notion of unicity distance by leveraging a philosophical interpretation of encryption as channel transmission. In this interpretation, the plaintext is seen as the message to be transmitted through a discrete memoryless channel (DMC) \cite{cover2006} (page 193)
to the attacker and the encryption is seen as the noise variable in this DMC model. The attacker is seen as attempting to decode the message transmitted over the DMC.

By utilizing the Channel Coding Theorem \cite{cover2006} (Theorem 7.7.1), we propose a new interpretation of unicity distance, i.e. the amount of text needed for communication to be reliable. This perspective allows us to analyze the behavior of a substitution cipher using reliable communication theory \cite{shannon1948}, ultimately providing an alternative proof of the unicity distance formula identical to Shannon's classical result. 

\section{Background: Entropy and Redundancy of Language}

Entropy of a random variable $X$ in the context of information theory, denoted as $H(X)$, measures the uncertainty or the information content in a random variable \cite{shannon1951}. For a language like English, entropy refers to the average information contained in each character or sequence of characters. The experimental work of Claude Shannon \cite{shannon1951} laid the foundation for understanding the entropy of printed English.

Redundancy refers to non-essential parts of the message that encode additional information. For English, redundancy is observed in linguistic conventions, such as common letter pairings or grammar rules. These conventions reduce the language’s entropy, making sequences of characters more predictable. For English text, the redundancy \( D \) is approximately 3.2 bits per letter \cite{hellman1977}.

\subsection{Classical Result: Unicity Distance}
A \textit{spurious key} is an incorrect key that also yields a meaningful decryption. For example, if the plaintext is \textit{cat} and it is encrypted into ciphertext $c$, then the key that deciphers $c$ into \textit{dog} would be a spurious key, since \textit{dog} is also a valid English word but it is not the actual plaintext. Intuitively, if the message is ``too short," then the number of spurious keys is inevitably nonzero, and the question we ask is how short is ``too short." To address this issue, we study the concept of unicity distance, which was first introduced by Shannon \cite{shannon1949}. The \textit{unicity distance} is the amount of ciphertext required on average for the number spurious key to be equal to $0$. To compute the unicity distance, the classical results \cite{shannon1949} and \cite{hellman1977} made the following assumptions:

\begin{enumerate}
    \item There are $2^{R_0 N}$ possible messages of length $N$. $R_0 = \log_2 G$ where $G$ is the number of letters in the alphabet of the Language. $G=26$ for English. 
    \begin{remark}
        For a message variable $(X_1,X_2,\dots,X_N)$ of length $N$. \[
        H(X_1,\dots,X_N)\leq \sum_{i=1}^N H(X_i)
        \] by the independence bound \cite{cover2006} (Theorem 2.6.6). For a language to achieve the maximum coding rate, the letters need to be mutually independent and uniformly distributed over the alphabet, i.e. $H(X_i)=\log_2 G=R_0$ \cite{cover2006}(Theorem 2.6.4). Therefore, $R_0$ represents the absolute rate of the language.
    \end{remark}
    \begin{remark}
        Assuming the letters are i.i.d., by the asymptotic equipartition property, almost all messages are typical and the number of typical sequences goes to $2^{R_0N}$ as $N\to \infty$.
    \end{remark}
    \item The number of meaningful messages of length $N$ is $2^{RN}$ distributed uniformly, where $R=\frac{H(\text{message}_N)}{N}$ is the entropy of the source message variable per letter. It is assumed that the rest of messages have negligible probabilities.
    \item There are $K$ keys which are a priori distributed uniformly and independent of the message variable.
\end{enumerate}
Shannon computed the unicity distance by considering the equivocation, the uncertainty of the key given a certain amount of intercepted ciphertext, and a subtle application of binomial distribution, without introducing the concept of spurious key \cite{shannon1949}. Hellman considered the number of spurious keys instead and the expected number of spurious keys \( \hat{n_k} \) is given by \cite{hellman1977}
\[
\hat{n_k} = (2^{H(K)}-1)2^{-ND}\approx 2^{H(K)-ND},
\] by arguing that the probability $p$ that randomly chosen key is spurious is given by $p=\frac{2^{RN}}{2^{R_0N}}$. This formula holds in a more general context as \cite{beauchemin1988} generalized Hellman's result so that there is actually no restriction on the distribution of the plaintext and the key. The unicity distance, $U$, as derived by Shannon and Hellman,  can be approximated using the following formula: 
\[
U = \frac{H(K)}{D}
\]
where $H(K)$ is the entropy of the key, and $D = R-R_0$ is the redundancy of the language. 
\begin{remark}
    In the context of unicity distance, it is assumed that a brute force attack is possible, as the information-theoretic framework does not concern itself with the specific nature of the attack. The classical unicity distance formula for English estimates that approximately 30 characters are required for an attacker to uniquely determine the key. More sophisticated statistical methods, such as the Metropolis-Hastings algorithm, typically require significantly longer plaintext to be effective, often requiring several hundred characters \cite{fathi2016}. However, as highlighted in the work of Diaconis \cite{diaconis2009}, Markov Chain Monte Carlo techniques successfully broke a substitution cipher with as few as 76 characters. Despite this success, there remains a considerable gap between the practical attack capabilities and the theoretical unicity distance, underscoring the limitations of advanced cryptanalytic techniques in achieving the theoretical minimum ciphertext length for unique key determination.
\end{remark}

\section{New Proof of the Classical Result}
\begin{proof}
We now fit the problem in the framework of channel coding theorem. Consider the DMC where the generic discrete channel $p(y|x)$ is defined as follows. $p(y|x)$ is a discrete channel where the message variable or the plaintext is the input random variable $X$, the enciphering transformation is the noise variable $Z$, and the ciphertext is the output random variable of the channel $Y$. This is indeed a discrete channel because the output $Y$ is deterministic function of $X$ and $Z$. The DMC we consider is just a sequence of independent replicates of this channel.

We now compute channel capacity of the DMC. Given the input distribution, which is the uniform distribution in our case, it suffices to consider the mutual information $I$ between the input and the output variables of the channel. There are two ways to apply the chain rule \cite{cover2006} (Theorem 2.4.1) to decompose the mutual information. The first way is as follows. 
\begin{eqnarray}
    I(\text{plaintext}_N, \text{ciphertext}_N) & = & H(\text{plaintext}_N) - H(\text{plaintext}_N | \text{ciphertext}_N)\\
    & = & \log_2 (2^{NR}) - \log_2(\#\text{spurious keys} + 1) \label{1}\\
    & = & NR - (H(K)-ND)\\
    & = & NR_0 - H(K)
\end{eqnarray}
Equality \ref{1} holds because given the ciphertext, the plaintext variable is distributed according to the keys yielding meaningful decipherment. This approach does not require any assumption on the plaintext distribution and the key distribution, by the generalization of Hellman's formula for the number of spurious keys in \cite{beauchemin1988}.

Alternatively, the mutual information can be decomposed as
\begin{eqnarray}
    I(\text{plaintext}_N, \text{ciphertext}_N) & = & H(\text{ciphertext}_N) - H(\text{ciphertext}_N | \text{plaintext}_N)\\
    & = & NR_0 - H(K) \label{2}
\end{eqnarray}
The second term in \ref{2} is $H(K)$ because given the plaintext, the ciphertext is distributed according to the chosen key, which is uniformly distributed. The claim that the first term in \ref{2} is equal to $NR_0$ may be a little far-fetched but in \cite{shannon1949}, Shannon specifically made the assumption that there are $2^{R_0}$ cryptograms (ciphertexts) and we assume that all ciphertexts are equally likely.
\begin{remark}
    \item The mutual information computed as above increases linearly with message length. In Shannon's approach, the equivocation decreases linearly with the message length and then follows an exponential decay \cite{shannon1949}. Also, note that the mutual information between the plaintext and the ciphertext is independent of the structure and redundancy of the language and is only dependent on the absolute rate of the language, i.e. the size of the alphabet.
\end{remark}
\begin{remark}
    The formula of the mutual information would yield a negative value if $N < \frac{H(K)}{R_0}$. In that case, we assume the mutual information is negligible since the substitution cipher behaves like a one-time pad for small message lengths.
\end{remark}

By the Channel Coding Theorem, there exists a code that allow the probability of error at the receiver of the DMC to be arbitrarily small if and only if the transmission rate $R_C$ is upper bounded by the channel capacity. We are interested in the case where the transmission rate is equal to the rate of the language $NR$. Therefore, we have reliable communication if and only if
\[
NR \leq NR_0 - H(K)
\]
Simplify to get
\[
\frac{H(K)}{D}\leq N.
\]
This yields the same unicity distance as Shannon’s classical result: \( \frac{H(K)}{D} \). In this context, the unicity distance can be viewed as the length of the message required on average for a substitution cipher secrecy system to become a reliable noisy communication system. 
\end{proof}

\section{Conclusion}

In this paper, we revisited the classical notion of unicity distance by framing encryption as a noisy transmission channel. By applying the Channel Coding Theorem, we derived a new proof of the unicity distance formula that is consistent with Shannon's original result. This channel-based interpretation not only reinforces the classical view but also provides further insight into the interpretation of unicity distance.

\section*{Acknowledgments}
The author acknowledges the guidance and support of Professor Steven N. Evans in shaping this work.

{\footnotesize  
\medskip
\medskip
\vspace*{1mm} 
 
\noindent {\it Fangyuan Lin}\\  
University of California, Berkeley\\
Berkeley, CA 94720, USA\\
E-mail: {\tt fangyuan@berkeley.edu}}
\end{document}